Sphere with mass 1 kg moving at speed v1 = 5m / s hits another sphere moving at speed

V2 = 2m / s face to face with the first. After the impact both spheres move together at a speed of 1m / s, according to the direction of the velocity of the second sphere before the impact. Find the mass of the second sphere.


1
Expert's answer
2021-02-17T10:34:18-0500

"m_1v_1-m_2v_2=(m_1+m_2)v\to"


"1\cdot 5-2m_2=1\cdot(1+m_2)\to m_2=4/3\ (kg)" . Answer